OpenStudy (jojomojo):

2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=?? Solve it?

jimthompson5910 (jim_thompson5910):

2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8= can be thought of as 2*(2+3-1) = 8 3*(3+7-1) = 27 4*(4+5-1) = 32 5*(5+8-1) = 60 6*(6+7-1) = 72 7*(7+8-1) =

jimthompson5910 (jim_thompson5910):

The general rule/pattern is x+y = x*(x+y-1)

OpenStudy (anonymous):

\[x+y=x[y+(x-1)]=x^2+xy-x\]\[7^2+7*8-7\]